Suppose a rope is to be strung form the top of two post of hights of 10 and 8 feet that are 24 feet apart. The rope needs to be staked to the ground between the two post. What is the shortest rope that can possible be used? How far from the taller post should it be staked?
Answer by checkley71(8403)

the solution I found(?) is 24/(4+5) 4+5 is the ratio of the 2 poles or
24/9=2 2/3 multiply 4*2 2/3=10.6666 & 5*2 2/3=13.3333 thus the hypotenuses of these two triangles are
A=SQRT10*10+13.333*13.333 OR A=SQRT100+177.768 OR A=SQRT277.768 OR A=16.6666 &
B=SQRT8*8+10.6666*10.6666 OR B=SQRT64+113.76 OR B=SQRT177.76 OR B=13.3333 THUS THE SUM OF THE HYPOTENUSES ARE 16.6666+13.3333 OR 30 FEET.
(I CHECKED MY ANSWER WITH A STRING AND 2 PINS)